# A faster algorithm for Vertex Cover parameterized by solution size

### v6

@@ 6 A simple branching algorithm
We now describe our first branching algorithm for vertex cover. As discussed earlier, the measure is a piecewise-linear function of $k$ and $\mu$. It is more convenient to describe it in terms of multiple self-contained “mini-algorithms,” with linear measure functions. As a starting point, we can describe the first algorithm:
Simplify $G$. If $\operatorname{maxdeg}(G)\leq 3$, then run either the algorithm of Theorem 2 or the MaxIS-3 algorithm, whichever is cheaper, and return. Otherwise, apply Section 5 to an arbitrary vertex of degree at least $4$, and run Branch4Simple on the two resulting subproblems.
Algorithm 1 Function Branch4Simple$(G,k)$
@@
Simplifying the graph in Line 1 can only reduce $\phi(G)=a\mu+bk$. If Section 5 is used, it gives a branch-seq $B$ with drops dominated by $[(1,3),(1,5)]$ or $[(0.5,1),(1.5,4)]$. To satisfy Eq. (3), we thus need:
@@
Although Proposition 27 is easy to check directly, the choices for the parameters $a$ and $b$ may seem mysterious. The explanation is that the most constraining part of the algorithm is dealing with the lower-degree graphs, i.e. solving the problem when the graph has maximum degree $3$. The inequality (5), which governs this case, should be completely tight: there should be a “triple point” with respect to Branch4Simple, the MaxIS-3 algorithm, and the algorithm of Theorem 2. That is, for chosen parameters $a,b$, we should have
$\mu\log(2.3146)=2(k-\mu)\log(1.083506)=a\mu+bk$
This allows us to determine $b$ in terms of $a$. Then our goal is to minimize $a$ whilst respecting the branching constraints of Eq. (4). This same reasoning will be used for parameters in all our algorithms. For example, we can define algorithms for branching on degree 5 and degree 6 vertices.
Simplify $G$ If $\operatorname{maxdeg}(G)\leq 4$, then run either the MaxIS-4 algorithm or Branch4Simple Otherwise, apply Section 5 to an arbitrary vertex of degree at least $5$.
Algorithm 2 Function Branch5Simple$(G,k)$
Simplify $G$ If $\operatorname{maxdeg}(G)\leq 5$, then run either the MaxIS-5 algorithm or Branch5Simple Otherwise, apply Section 5 to an arbitrary vertex of degree at least $6$.
Algorithm 3 Function Branch6Simple$(G,k)$
Along similar lines, we immediately obtain the results:
@@
Combined with the MaxIS-5 algorithm using Lemma 10, Branch5Simple immediately gives an algorithm for degree-5 graphs with runtime $O^{*}(1.2491^{k})$. The final algorithm is very similar, but we only keep track of runtime in terms of $k$, not $\mu$ or $n$.
If $\operatorname{maxdeg}(G)\leq 6$, then use Lemma 10 with algorithms of Branch6Simple and MaxIS-6 Otherwise, split on an arbitrary vertex of degree at least $7$.
Algorithm 4 Function Branch7Simple$(G,k)$
@@
Algorithm Branch7Simple$(G,k)$ runs in time $O^{*}(1.2575^{k})$.
@@
If we split on a vertex of degree at least $7$, then we generate subproblems with vertex covers of size at most $k-1$ and $k-7$. These have cost $1.2575^{k-1}$ and $1.2575^{k-7}$ by induction. Since $11.2575^{-1}+1.2575^{-7}<$ 1, the overall cost is $O^{*}(1.2575^{k})$. Otherwise the two algorithms in question have measure $a\mu+bk$ and $cn$ respectively; where $a=0.20199,b=0.160637,c=\log(1.1893)$; by applying Lemma 10, we get a combined algorithm with cost $O^{*}(1.2575^{k})$ as desired. ∎
@@ 7 The advanced branching algorithm: overview
In the next sections, we consider more elaborate graph structures and branching rules. This improves the runtime to $O^{*}(1.25284^{k})$, along with improved algorithms for lower-degree graphs. The outline is similar to Section 6: we develop a series of branching algorithms targeted to graphs with different degrees, building our way up from degree-4 graphs to degree-7 graphs. Here we highlight some of the main changes in our analysis.
Multiway branching. So far, we have only analyzed two-way branching rules. The advanced algorithms will branch on multiple vertices at a time. Specifically, we say that we split on a sequence of vertices $x_{1},\dots,x_{\ell}$ if we first split on $x_{1}$; in the subproblem $\langle G-x_{1},k-1\rangle$ we split on $x_{2}$; in the subproblem $\langle G-\{x_{1},x_{2}\},k-2\rangle$ we split on $x_{3}$; and so on. Assuming that $x_{1}\dots,x_{\ell}$ are independent, this generates $\ell+1$ subproblems $\langle G-N[x_{1}],k-\deg(x_{1})\rangle,\langle G-x_{1}-N[x_{2}],k-1-\deg(x_{2})\rangle,\langle G-\{x_{1},x_{2}\}-N[x_{3}],k-2-\deg(x_{3})\rangle,\dots,\langle G-\{x_{1},\dots,x_{\ell-1}\}-N[x_{\ell}],k-(\ell-1)-\deg(x_{\ell})\rangle,\langle G-\{x_{1},\dots,x_{\ell}\},k-\ell\rangle$. We emphasize that the subproblem $G-N[x_{1}]$ does not have any further branching, such as guessing the status of $x_{2}$. Splitting on $x_{1},x_{2}$ is just a 3-way branch.
Miscellaneous branching rules. We occasionally use some other branching rules; some of these can be applied to graphs which are not simplified. The following rules are used repeatedly:
(A1): Branch on 2-vertex. For a 2-vertex $v$ with neighbors $x$, $y$, Lemma 5 implies there is either a good cover omitting $v$, or omitting both $x,y$. Thus, we can branch on subproblems $\langle G-N[v],k-|N(v)|\rangle$ and $\langle G-N[x,y],k-|N(x,y)|\rangle$.
(A2): Branch on a funnel. For a funnel $u$ with out-neighbor $x$ and any other neighbor $v$ of $u$, we branch on whether $v\in C$ or whether $\{v,x\}\cap C=\emptyset$; this generates subproblems $\langle G-v,k-1\rangle$ and $\langle G-N[v,x],k-|N(v,x)|\rangle$.
(A3): Branch on shared neighbors. For vertices $u,v$ branch on whether $\{u,v\}\subseteq C$, or $N(u)\cap N(v)\subseteq C$; this generates subproblems $\langle G-\{u,v\},k-2\rangle$ and $\langle G-N(u)\cap N(v),k-\operatorname{codeg}(u,v)\rangle$.
We use a few specialized branching rules in various places; these will be discussed as needed.
Creation of high-degree vertices and surplus-two indsets. One common tactic in our analysis is to branch in a targeted way to allow us the use a particularly profitable “follow-up” branching rule in the resulting graph. Two of the most common situations are to use preprocessing rules P2 or P3 in order to create new high-degree vertices, or branching to create a large surplus-two indset for which we can use brancing rules in Section 5.1.
We say graph $G$ is $r$-vulnerable (or $r$-vul for short) if either (i) $S(G)\geq 1$ or (ii) $G$ has an indset with surplus at most two and size at least $r$.
@@
In either case, let $r=\deg(x)+\deg(y)-\operatorname{codeg}(x,y)-1$.
If $r\geq 5$, then $G$ has available branch-seq $[(0,2)]$, $[(1,4),(1,6)]$ or $[(0.5,2),(2,6)]$. If $r\geq 6$ then $G$ has available branch-seq $[(0,2)]$, $[(1,4),(1,6)]$, $[(0.5,3),(2,6)]$, or $[(0.5,2),(2.5,r+1)]$.
@@
In case (i), we apply (P2) to merge $x,y$ into a new $r$-vertex $z$. In case (ii), we apply (P3); if $\operatorname{codeg}(u,x)\geq 0$ this directly gives drop $[(0,2)]$. Otherwise, it produces an $r$-vertex $y$. If the resulting graph $G^{\prime}$ can be simplified further (aside from isolated vertices), then we get net drop $(0,2)$. Otherwise, we apply Section 5 to $G^{\prime}$. ∎
@@
Suppose $\operatorname{mindeg}(G)\geq 1$.
@@
If $G$ is not simplified, then $S(G)\geq 1$ giving branch-seq $(0,1)$. Otherwise, let $I$ be a surplus-two min-set of size $r$. If $r\geq 4$ then splitting on $I$ gives subproblems $G-I,G-N[I]$ with drops $(1,r)$ and $(1,r+2)$ respectively. If $r=2$ or $r=3$, we apply Lemma 22 or Lemma 21 respectively. ∎
For Propositions 30 and 31, we use the following notation for $\operatorname{val}_{a,b}(B)$ of the given branch-seq $B$.
@@
Notational convention: omitting the value $k$. For the remainder of this paper, when analyzing branching subproblems, we will usually omit the solution size $k^{\prime}$. It should be determined from context, from the following mnemonic: for a subproblem $G-\{x_{1},\dots,x_{\ell}\}-N[y]$, we take $k^{\prime}=k-\ell-\deg(y)$. In using this notation, we must be careful to ensure that $\{x_{1},\dots,x_{\ell}\}\not\sim y$, as otherwise some $x_{i}$ may be double-counted.
@@ 8 Branching on 4-vertices
We now develop an improved algorithm Branch4. At a high level, the improvement over the baseline algorithm Branch4Simple comes from tracking the presence of 3-vertices. If a 4-vertex $u$ has a 3-neighbor $v$, then we can split on $u$ and then, in the branch $G-u$, the degree of $v$ gets reduced to two. We then use (P2) to reduce $k$ by one – a “bonus” reduction. We can also ensure that this creates new 3-vertices; thus, the process is a self-sustaining chain reaction.
@@
Apply branching rules to ensure $G$ has nice structure, e.g. simplified and no 4-cycles.                  (We will call this 4-base-position (4BP) later.) If $V_{3}\neq\emptyset$ and $V_{3}\not\sim V_{4}$, then solve vertex cover exactly on $V_{3}$. Split on a 4-vertex with a 3-neighbor, chosen according to tie-breaking rules to be described later. Branch on a triangle. Branch on a vertex $u$ according to tie-breaking rule to be described later.
Algorithm 5 Function Branch4$(G,k)$
@@
Via Lemma 10, in combination with the MaxIS-4 algorithm, this will give our algorithm with runtime $O^{*}(1.21131^{k})$ for degree-4 graphs. Theorem 32 is very length to prove. One major complication is that it requires us to use a measure for $G$ tracking additional information beyond just the values $\mu$ and $k$. Specifically, our strategy is to show that Branch4 has the measure of the form
$\phi(G)=a\mu+bk-\operatorname{bonus}(G);$
where we define function $\operatorname{bonus}(G)$ as follows. First, we define a special 4-vertex to be a 4-vertex with at least two 3-neighbors, and we say $G$ is special if it has a special 4-vertex. We then define:
$\operatorname{bonus}(G)$ =$=\begin{cases}\beta&\text{if $\operatorname{mindeg}(G)=3$ and $\operatorname{maxdeg}(G)=4$ and $G$ is special}\\ \alpha&\text{if $G$ has a subcubic vertex and the previous case does not hold}\\ 0&\text{otherwise}\end{cases}$
@@
Motivated by Eq. (6), we write $S(G)\geq\ell^{*}$ if either (i) $S(G)\geq\ell+1$, or (ii) $S(G)=\ell$ and there is a simplified graph $\langle G^{\prime},k-\ell\rangle$ with a 3-vertex. We also write $R(X)\geq\ell^{*}$ if $S(G-X)\geq\ell^{*}$, etc.
Similarly, if we transform a graph $\langle G,k\rangle$ into a graph $\langle G^{\prime},k^{\prime}\rangle$, we say $G^{\prime}$ has “extended drop” $(\Delta\mu,\Delta k^{*})$ if $\mu^{\prime}\leq\mu-\Delta\mu$ and either (i) $G^{\prime}$ has a subcubic vertex and $k^{\prime}=k-\Delta k$ or (ii) $k^{\prime}<k-\Delta k$ strictly. To explain, note that since $\alpha\leq b$, we have $\phi(G^{\prime})\leq\phi(G)-a\Delta\mu-b\Delta k-\alpha+\operatorname{bonus}(G)$. So the contribution of $G^{\prime}$ to the sum in Eq. (3) is at most $e^{-a\Delta\mu-b\Delta k-\alpha}$.
8.1 Some observations about simplification and degree-3 vertices
We record a few general observations about simplification in graphs.
Proposition 33.
Let $s\geq 1$. If $G$ has at least $2(s+1)$ subcubic vertices, and it has a critical-set $I$ with $\operatorname{surp}(I)=1,|I|\geq s$, then $S(G)\geq s^{*}$.
Proof.
We apply (P2) to $I$. If $|I|>s$, this gives $S(G)\geq s+1$ directly. Otherwise, all subcubic vertices outside $N[I]$ remain subcubic, leaving at least $12(s+1)-|N[I]|=$ 1 subcubic vertices.
Proposition 34.
Suppose $G$ has no 4-cycles, has at least 7 subcubic vertices, and has two non-adjacent 2-vertices. Then $S(G)\geq 2^{*}$.
Proof.
Let $u,v$ be non-adjacent 2-vertices. If we apply (P2) to $v$, then all the subcubic vertices in $G$ outside from $N[v]$ remain subcubic in the resulting graph $G^{\prime}$. Since $G$ has no 4-cycle, we have $\deg_{G^{\prime}}(u)=2$. So $G^{\prime}$ has at least 4 subcubic vertices and has a 2-vertex $u$. By Proposition 33, this implies $S(G^{\prime})\geq 1^{*}$ and hence $S(G)\geq 2^{*}$. ∎
Lemma 35.
Let $X$ be a vertex set. If $\operatorname{maxdeg}(G)\leq 4$ and $G$ has no components of size at most $8|X|$, then $G-X$ has a subcubic vertex.
Proof.
There must be some vertex $z$ at distance two from $X$, since otherwise the component(s) containing $X$ would have size at most $(4+\operatorname{maxdeg}(G))|X|\leq 8|X|$. So suppose $z$ has a neighbor $y\sim X$. Here $\deg_{G-X}(y)\geq 1$ (due to neighbor $z$) and also $\deg_{G-X}(y)\leq 3$ (it loses some neighbor in $X$). Hence, $G-X$ has a subcubic vertex. ∎
Proposition 36.
Let $r\geq 2$ be an arbitrary constant. Suppose that $G$ has at least 4 subcubic vertices and $\operatorname{maxdeg}(G)\leq 4$ and $\operatorname{minsurp}(G)\geq 1$ and $G$ has an indset $I$ with $|I|\geq r$ and $\operatorname{surp}_{G}(I)\leq 2$. Then $G$ has available branch-seq $[(0,1^{*})]$ or $[(1,r^{*}),(1,(r+2)^{*})]$.
Proof.
If $G$ has any components of size less than $8(r+2)$, we solve them exactly in $O(1)$ time and remove them from $G$; since $\operatorname{minsurp}(G)\geq 1$, this reduces $k$ by at least two and hence gives drop $[(0,2)]$. If $\operatorname{surp}_{G}(I)=1$, then Proposition 33 with $s=1$ gives $S(G)\geq 1^{*}$. Otherwise, we split on min-set $I$; if $|I|>r$ strictly, the subproblem $G-I$ and $G-N[I]$ have drop $(1,(r+1))$ and $(1,(r+3))$ directly. If $|I|=r$, these subproblems have drops $(1,r)$ and $(1,r+2)$ and, by Lemma 35, both of them have subcubic vertices.
We say that $G$ is $r^{*}$-vul if either $S(G)\geq 1^{*}$ or the preconditions of Proposition 36 hold, and we correspondingly define
$\gamma^{*}_{r}=\max\{e^{-b-\alpha},e^{-a-rb-\alpha}+e^{-a-(r+2)b-\alpha}\}$

### v7

@@ 4 A simple branching algorithm
We now describe our first branching algorithm for vertex cover, where the measure is a piecewise-linear function of $k$ and $\mu$. It is more convenient to describe this in terms of multiple self-contained “mini-algorithms,” with linear measure functions.
As a starting point, we can describe the first mini-algorithm:
1Simplify $G$.
2If $\operatorname{maxdeg}(G)\leq 3$, then run either the algorithm of Theorem 1 or the MaxIS-3 algorithm, whichever is cheaper, and return.
Otherwise, apply Figure 6 to an arbitrary vertex of degree at least $4$, and run Branch4Simple on the two resulting subproblems.
Algorithm 1 Function Branch4Simple(G,k)(G,k)
@@
Simplifying the graph in Line 1 can only reduce $\phi(G)=a\mu+bk$. If Figure 6 is used, it gives a branch-seq $B$ with drops dominated by $[(1,3),(1,5)]$ or $[(0.5,1),(1.5,4)]$. To satisfy Eq. (2), we thus need:
@@
Although Proposition 27 is easy to check directly, the choices for the parameters $a$ and $b$ may seem mysterious. The explanation is that the most constraining part of the algorithm is dealing with the lower-degree graphs, i.e. solving the problem when the graph has maximum degree $3$. The inequality (4), which governs this case, should exhibit a “triple point” with respect to Branch4Simple, the MaxIS-3 algorithm, and the algorithm of Theorem 1. That is, for chosen parameters $a,b$, we should have
$\mu\log(2.3146)=2(k-\mu)\log(1.083506)=a\mu+bk.$
This allows us to determine $b$ in terms of $a$. Then our goal is to minimize $a$ whilst respecting the branching constraints of Eq. (3). This same reasoning will be used for parameters in all our algorithms.
Likewise, we define algorithms for branching on degree 5 and degree 6 vertices.
1Simplify $G$
2If $\operatorname{maxdeg}(G)\leq 4$, then run either the MaxIS-4 algorithm or Branch4Simple
Otherwise, apply Figure 6 to an arbitrary vertex of degree at least $5$.
Algorithm 2 Function Branch5Simple(G,k)(G,k)
1Simplify $G$
2If $\operatorname{maxdeg}(G)\leq 5$, then run either the MaxIS-5 algorithm or Branch5Simple
Otherwise, apply Figure 6 to an arbitrary vertex of degree at least $6$.
Algorithm 3 Function Branch6Simple(G,k)(G,k)
@@
The final algorithm is very similar, but we only keep track of runtime in terms of $k$, not $\mu$ or $n$.
1If $\operatorname{maxdeg}(G)\leq 6$, then use Lemma 15 with algorithms of Branch6Simple and MaxIS-6
Otherwise, split on an arbitrary vertex of degree at least $7$.
Algorithm 4 Function Branch7Simple(G,k)(G,k)
@@
Algorithm Branch7Simple(G,k)(G,k) runs in time $O^{*}(1.2575^{k})$.
@@
If we split on a vertex of degree at least $7$, then we generate subproblems with vertex covers of size at most k1k-1 and k7k-7. These have cost $1.2575^{k-1}$ and $1.2575^{k-7}$ by induction. Since $1.2575^{-1}+1.2575^{-7}<1$, the overall cost is $O^{*}(1.2575^{k})$. Otherwise the two algorithms in question have measure $a\mu+bk$ and $cn$ respectively; where $a=0.20199,b=0.160637,c=\log(1.1893)$; by applying Lemma 15, we get a combined algorithm with cost $O^{*}(1.2575^{k})$ as desired. ∎
@@ 5 The advanced branching algorithm: overview
In the next sections, we consider more elaborate graph structures and branching rules. This improves the runtime to $O^{*}(1.25284^{k})$, along with algorithms for lower-degree graphs. The outline is similar to Section 4: we develop a series of branching algorithms targeted to graphs with different degrees, building our way up from degree-4 graphs to degree-7 graphs. Here we highlight some of the main changes in our analysis.
5.1 Multiway branching
So far, we have only analyzed two-way branching rules. The advanced algorithms will use multiway branching; for example, we may split on a given vertex $u$, and then apply further branching steps to the subproblems GuG-u and/or GN[u]G-N[u].
A common tactic, which we refer to as splitting on a sequence of vertices $x_{1},\dots,x_{\ell}$, is to first split on $x_{1}$; in the subproblem $\langle G-x_{1},k-1\rangle$ then split on $x_{2}$; in the subproblem $\langle G-\{x_{1},x_{2}\},k-2\rangle$ then split on $x_{3}$, and so on. When $x_{1}\dots,x_{\ell}$ are independent, this generates subproblems $G_{1}=\langle G-N[x_{1}],k-\deg(x_{1})\rangle,G_{2}=\langle G-x_{1}-N[x_{2}],k-1-\deg(x_{2})\rangle,\dots,G_{\ell}=\langle G-\{x_{1},\dots,x_{\ell-1}\}-N[x_{\ell}],k-\ell+1-\deg(x_{\ell})\rangle,G_{0}=\langle G-\{x_{1},\dots,x_{\ell}\},k$-$\ell\rangle$. We emphasize that $G_{1}$ does not have further branching, such as guessing the status of $x_{2}$. Splitting on $x_{1},x_{2}$ is just a 3-way branch.
A related tactic is to branch in a targeted way to create large surplus-two indsets, or so that preprocessing rules (P2) or (P3) create new high-degree vertices. The branching rules in Section 3.1 can then be applied, which are highly profitable. We record a few examples of this analysis here.
@@ 5.1 Multiway branching
In either case, define $r=\deg(x)+\deg(y)-\operatorname{codeg}(x,y)-1.$
If $r\geq 5$ then $G$ has available branch-seq $[(0,2)]$, $[(1,4),(1,6)]$ or $[(0.5,2),(2,6)]$.
If $r\geq 6$ then $G$ has available branch-seq $[(0,2)]$, $[(1,4),(1,6)]$, $[(0.5,3),(2,6)]$, or $[(0.5,2),(2.5,r+1)]$.
@@
In case (i), we apply (P2) to merge $x,y$ into a new $r$-vertex $z$. In case (ii), we apply (P3); if $\operatorname{codeg}(u,x)\geq 0$ this directly gives drop $(0,2)$. Otherwise, it produces an $r$-vertex $y$. If the resulting graph $G^{\prime}$ can be simplified further (aside from isolated vertices), then we get net drop $(0,2)$. Otherwise, we apply Figure 6 to $G^{\prime}$. ∎
@@
Suppose $\operatorname{mindeg}(G)\geq 1$. We say graph $G$ is $r$-vulnerable (or $r$-vul for short) if either (i) $S(G)\geq 1$ or (ii) $G$ has an indset with surplus at most two and size at least $r$.
@@
If $G$ is not simplified, then $S(G)\geq 1$ giving branch-seq $(0,1)$. Otherwise, we can apply Lemma 22 or Lemma 21 or just apply (B-Crit) depending on the size of the indset. ∎
Notation for branching rules. For the remainder of this paper, when analyzing branching subproblems, we usually omit the solution size $k^{\prime}$. It should be determined from context: a subproblem $G^{\prime}=G-X-N[Y]$ is principal with excess $|Y|$ and has $k^{\prime}=k-|X|-|N(Y)|$. (For this notation, we are careful to ensure that $Y$ is an indset and that $X\not\sim Y$, to avoid double-counting.)
We use the following important compositional property: if we generate subproblems $i=1,\dots,t$ each with drop $(\Delta\mu_{i},\Delta k_{i})$, and then apply a branching rule with branch-seq $B^{\prime}_{i}$ to the subproblem $i$, then the resulting branching rule $B$ overall has
$\operatorname{val}_{a,b}(B)=\sum_{i=1}^{t}e^{-a\Delta\mu_{i}-b\Delta k_{i}}\operatorname{val}_{a,b}(B_{i}^{\prime}).$
For Propositions 30 and 31, we use the following notation for the $\operatorname{val}(B^{\prime}_{i})$:
@@
Thus, for instance, if we have a subproblem with drop $(\Delta\mu,\Delta k)$ and then we apply Proposition 30, its overall contribution to the Branching Inequality, would be given by $e^{-a\Delta\mu-b\Delta k}\psi_{r}.$
5.2 Branching on funnels and other structures
If the graph contains a funnel, the simplest tactic is to apply (P3). However this creates additional edges, which can make the graph too complicated to analyze further. It is often better to branch on the funnel itself. There are two possible branching rules for this situation:
(B-Fun1) For a funnel $u$ with out-neighbor $x$, branch on whether $u\notin C$ or $x\notin C$, generating subproblems GN[u]G-N[u] and GN[x]G-N[x]. [^7] Recall our notational convention that these subproblems have $k^{\prime}=k-\deg(u)$ and $k^{\prime}=k-\deg(x)$ respectively.
(B-Fun2) For a funnel $u$ with out-neighbor $x$ and any other neighbor $v$ of $u$, branch on whether $v\in C$ or whether $\{v,x\}\cap C=\emptyset$; this generates subproblems GvG-v and GN[v,x]G-N[v,x].
Note that, when we apply (B-Fun2), the vertex $u$ still has a funnel in the subproblem GvG-v (of size reduced by one); we can continue to apply (P3) to it, or use other funnel reductions.
Consider a degree-two vertex $u$ with neighbors $x,y$. This can be regarded as a funnel, where $x$ is an out-neighbor and the vertex $\{y\}$ forms a (trivial) clique. If we apply (B-Fun2), we get subproblems GyG-y and GN[x,y]G-N[x,y]. In the former subproblem, $u$ now has degree one; we can apply (P1) to remove $u$ and its neighbor $x$ and get subproblem $G-\{x,y\}=G-N[u]$. Since this comes up so frequently, we summarize this branching strategy as follows:
(B-Pair) For a 2-vertex $u$ with neighbors $x,y$ branch on subproblems $G-\{x,y\}$ and GN[x,y]G-N[x,y].
We occasionally use some other specialized branching rules; most of these will be discussed in the specific situations where they are needed. There is one branching rule worth further discussion, as it comes up in a number of contexts:
(B-Shared) Suppose vertices $u,v$ share neighbors $x_{1},\dots,x_{\ell}$. Branch on whether $\{u,v\}\subseteq C$, or $\{x_{1},\dots,x_{\ell}\}\subseteq C$; this generates subproblems $G-\{u,v\}$ and $G-\{x_{1},\dots,x_{\ell}\}$.
Note that, in applying (B-Shared), it is allowed to have $\operatorname{codeg}(u,v)>\ell$ strictly.
5.3 Tracking degree-3 vertices
In the algorithms we have encountered so far, we only use two parameters of the graph: the values $k$ and $\mu$. This will also be the case for most of our advanced algorithms. For the degree-4 algorithm, however, we will need to track a few additional pieces of information. The most important is whether the graph has any degree-3 vertices (the precise number of them is not tracked). These would lead to more profitable branching rules; as a rule of thumb, having a degree-3 vertex is almost as good as simplifying the graph.
For the degree-4 algorithm, we extend our notations to track 3-vertices. We write $S(G)\geq\ell^{*}$ if either (i) $S(G)=\ell$ and the simplified graph $\langle G^{\prime},k$-$\ell\rangle$ has a 3-vertex, or (ii) $S(G)\geq\ell+1$. We say that a subproblem $\langle G^{\prime},k^{\prime}\rangle$ has drop $(\Delta\mu,\Delta k^{*})$ if $\mu^{\prime}\leq\mu-\Delta\mu$ and either (i) $G^{\prime}$ has a subcubic vertex and $k^{\prime}=k-\Delta k$ or (ii) $k^{\prime}<k-\Delta k$ strictly.
Here we record a few observations about simplification and degree-3 vertices.
Proposition 32.
Let $s\geq 1$. If $G$ has at least $2(s+1)$ subcubic vertices, and it has a critical-set $I$ with $\operatorname{surp}_{G}(I)=1,|I|\geq s$, then $S(G)\geq s^{*}$.
Proof.
Apply (P2) to $I$. If $|I|>s$, this gives $S(G)\geq s+1$ directly. Otherwise, all subcubic vertices outside $N[I]$ remain subcubic, leaving at least 2(s+1)|N[I]|=12(s+1)-|N[I]|=1 subcubic vertices. ∎
Proposition 33.
If $G$ has no 4-cycles, has at least 7 subcubic vertices, and has two non-adjacent 2-vertices, then $S(G)\geq 2^{*}$.
Proof.
Let $u,v$ be non-adjacent 2-vertices. Apply (P2) to $v$; then all the subcubic vertices in $G$ outside from $N[v]$ remain subcubic in the resulting graph $G^{\prime}$. So $G^{\prime}$ has at least 4 subcubic vertices. Also, since $G$ has no 4-cycle, we have $\deg_{G^{\prime}}(u)=2$. By Proposition 32, this implies $S(G^{\prime})\geq 1^{*}$ and hence $S(G)\geq 2^{*}$. ∎
Lemma 34.
Suppose that $\operatorname{maxdeg}(G)\leq 4$, and let $G^{\prime}=\langle G-(X\cup Y),k-|X|\rangle$ be a principal subproblem. If all components of $G$ have size at least $5|X|+1$, then $S(G^{\prime})$ has drop $(\Delta\mu,|X|^{*})$.
Proof.
There must be some vertex $z$ at distance two from $X$, since otherwise the component(s) containing $X$ would have size at most $5|X|$. So suppose $z$ has a neighbor $y\sim X$. Here $\deg_{G-X}(y)\geq 1$ (due to neighbor $z$) and also $\deg_{G-X}(y)\leq 3$ (it loses some neighbor in $X$). So GXG-X has a subcubic vertex. Removing isolated vertices does not affect this. ∎
Proposition 35.
Let $r\geq 2$ be an arbitrary constant. Suppose that $G$ has at least 4 subcubic vertices and $\operatorname{maxdeg}(G)\leq 4$ and $\operatorname{minsurp}(G)\geq 1$ and $G$ has an indset $I$ with $|I|\geq r$ and $\operatorname{surp}_{G}(I)\leq 2$. Then $G$ has available branch-seq $(0,1^{*})]$ or $(1,r^{*}),(1,(r+2)^{*})]$.
Proof.
If $G$ has any components of size at most $5(r+2)$, we solve them exactly in constant time and remove them from $G$; since $\operatorname{minsurp}(G)\geq 1$, this reduces $k$ by at least two and hence gives drop $[(0,2)]$. If $\operatorname{surp}_{G}(I)=1$, then Proposition 32 with $s=1$ gives $S(G)\geq 1^{*}$. Otherwise, we apply (B-Crit); by Lemma 34 the subproblems have drops $(1,r^{*})$ and $(1,(r+2)^{*})$. ∎
Along the lines of Proposition 31, we say that $G$ is $r^{*}$-vul if either $S(G)\geq 1^{*}$ or the preconditions of Proposition 35 hold.
@@ 6 Branching on 4-vertices
We now develop an improved algorithm Branch4. At a high level, the improvement over the baseline algorithm Branch4Simple comes from tracking the presence of 3-vertices. If a 4-vertex $u$ has a 3-neighbor $v$, then we can split on $u$ and then, in the branch GuG-u, the degree of $v$ gets reduced to two. We then use (P2) to reduce $k$ by one – a “bonus” reduction. We can also ensure that this creates new 3-vertices; thus, the process is a self-sustaining chain reaction.
@@
1
2Apply misc. branching rules to ensure $G$ has nice structure, e.g. simplified and no 4-cycles.
3If $V_{3}\neq\emptyset$ and $V_{3}\not\sim V_{4}$, then solve vertex cover exactly on $V_{3}$.
4Split on a 4-vertex with a 3-neighbor, chosen according to tie-breaking rules to be described later.
5Branch on a triangle.
Branch on a vertex $u$ according to tie-breaking rule to be described later.
Algorithm 5 Function Branch4(G,k)(G,k)
@@
Via Lemma 15, in combination with the MaxIS-4 algorithm, this will give our algorithm with runtime $O^{*}(1.21131^{k})$ for degree-4 graphs. Theorem 36 is very lengthy to prove. Our strategy is to show that Branch4 has the measure of the form
$\phi(G)=a\mu+bk-\operatorname{bonus}(G).$
Let us define a special 4-vertex to be a 4-vertex with at least two 3-neighbors, and we say $G$ is special if it has a special 4-vertex. We then define:
[…]
@@
To show that a branching rule generating subproblems $G_{1},\dots,G_{t}$ respects the measure $\phi$, we using a slight variant of the branching inequality:
$\sum_{i=1}^{t}e^{\phi(G_{i})-a\mu-bk}\leq e^{-\operatorname{bonus}(G)}$ (6)
As a consequence of Eq. (5) and Proposition 35, we have the following observation (which motivates our notation for tracking 3-vertices as we have described earlier):
Observation 37.
If $S(G)\geq\ell^{*}$, then the simplified graph $G^{\prime}$ has $\phi(G^{\prime})\leq\phi(G)-b\Delta k-\alpha$.
If subproblem has drop $(\Delta\mu,\Delta k^{*})$, then its contribution to Eq. (6) is at most $e^{-a\Delta\mu-b\Delta k-\alpha}$.
If subproblem has drop $(\Delta\mu,\Delta k)$ and is $r^{*}$-vul, its contribution to Eq. (6) is at most $e^{-a\Delta\mu-b\Delta k}\gamma^{*}_{r}$ where we define $\gamma^{*}_{r}=\max\{e^{-b-\alpha},e^{-a-rb-\alpha}+e^{-a-(r+2)b-\alpha}\}$.